In a circus the diameter of the globe of death is 20m. From what minimum height must a cyclist start in order to go round the globe vertically? Please provide solution

Grade:10

2 Answers

Manas Shukla
7 years ago
At highest point of globe of death , the energy equation will be
mg 2R + \frac{1}{2} m v^{2} = m g h
Where R is the radius of the globe of death.
We also know at highest point mototcyclist needs to cancel gravityto prevent from falling down.
Which will be provided by the centripetal acceleration.
thus \frac{v^{2}}{R} = g
v2 = Rg
substituting it in the energy equation and solving we get
h = \frac{5R}{2}
h =50m
Gosula Madhukar
7 years ago
hmin=radius/2=5m.
i hope iam correct.
plz give reply.
plz post such type of questions.
have a nice day.